OpenStudy (anonymous):

17. y = x2 + 3x – 4 y = 2x + 2

OpenStudy (anonymous):

what do u need to do with this problem?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

you have y = 2x + 2 plug that in for the initial equation then solve for x

OpenStudy (anonymous):

can you show me how

OpenStudy (anonymous):

2x + 2 = x2 + 3x - 4 x2 + 3x - 2x = 2 + 4 can you continue?

OpenStudy (anonymous):

no what else do to do

OpenStudy (anonymous):

x^2 + x - 6 = 0 factor it out you have (x+3)(x-2) = 0 solve for x then plug your answer in to find y

OpenStudy (anonymous):

what else do you do after this

OpenStudy (anonymous):

what are your answers for x?

OpenStudy (anonymous):

i don know hwo to do that

OpenStudy (anonymous):

x + 3 = 0 x - 2 = 0 solve for that

OpenStudy (anonymous):



OpenStudy (anonymous):

yep. then plug that in the second equation to find y

OpenStudy (anonymous):

ok so the answer is 8

OpenStudy (anonymous):

@ninhi5 answer is 8 or 6

OpenStudy (anonymous):

(–3, 6) and (2, –4) (–3, –4) and (2, 6) (–3, –4) and (–2, –2) no solution

OpenStudy (anonymous):

pick one :)

OpenStudy (anonymous):

the 1st or 2nd

OpenStudy (anonymous):

nope nope plug in -3 for y = 2x + 2 to solve for y => y = 2(-3) + 2 = ? then do the same for -2

OpenStudy (anonymous):

-4 for the 1st then -2 for the second

OpenStudy (anonymous):

IT IS THE THIRD

OpenStudy (anonymous):

Hooraaaayyyy! good job :)

OpenStudy (anonymous):

thankscan you help me with my other one